8 Total Question 3 A clothing company changes its sales depending on if the weather is clear or raining. A weather forecastin3(c). In the long run, what proportion of days should the company prepare 5 marks sales for clear weather?
Answer #1

(a)

Let the states be C and R denoting that it is clear or raining today respectively.
The transition probability from C to C is 0.8
The transition probability from C to R is 1 - 0.8 = 0.2
The transition probability from R to R is 0.65
The transition probability from R to C is 1 - 0.65 = 0.35


The transition matrix is,

P= 0.8 0.35 0.35 0.2 0.65

(b)

Probability that it will rain next Tuesday = Probability that it will rain on Monday * transition probability from R to R + Probability that it will be clear on Monday * transition probability from C to R

= (1 - 0.2) * 0.65 + 0.2 * 0.2

= 0.56

(c)

Let \pi = [a b] be the long run proportions of clear or raining weather

Then,  \piP = \pi and a + b = 1

which gives,

0.8a + 0.35b = a => 0.2a = 0.35b => b = 0.5714286 a

0.2a + 0.65b = b

a + b = 1 => a + 0.5714286 a = 1

=> 1.5714286 a = 1

=> a = 1/ 1.5714286 = 0.6364

Thus, the long run proportions of clear weather is 0.6364